OpenStudy (anonymous):

In January of the year 200, my wife Lisa was eleven times as old as my son Don. In January 2012, she will be three times as old as my son. How old was my son in January of 2000? ~ Anyone have the steps, equation, and answers... <3

OpenStudy (anonymous):

let x be the son's age in 2000 can you express lisa's age in 2000 in terms of x?

OpenStudy (anonymous):

idk how to do the problem?

OpenStudy (anonymous):

if don's age was x,his mother was 11 times as old as he was so she was 11x can you follow me so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good in 2012 don is 12 years older than in 2000 so he's now (x+12) the mother is also 12 years older so she's now (11x+12) good?

OpenStudy (anonymous):

soo its 11x+(11x+12)

OpenStudy (anonymous):

just wait a second ,did you understand my last posting?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good ,now we need an equation to solve for x they say ,in 2012 lisa is 3 times as old as don so her age now=3*(don's age now) =>(11x+12)=3(x+12) can you take it from here?

OpenStudy (anonymous):

yes thank you soo much :))

OpenStudy (anonymous):

good,you're welcome